Write the sequence of the following explicit formula.

1. sn = (-4)^n, 1 ≤ n

2. bn = 92 - 5n, 1 ≤ n ≤ 5

3. Blue taxi inc. charges $7.50 for the first 4 miles and 1 dollar for each additional miles. The table shows the cost of travelling from 4 to 10 miles.

mileage 4, cost 7.50

mileage 5, cost 8.50

mileage 6, cost 9.50

mileage 7, cost 10.50

mileage 8, cost 11. 50

mileage 9, cost 12.50

mileage 10, cost 13. 50


Write the recursive formula.

Write the explicit formula


4. If the sequence formula is given as tn = n^2 -1, n less than or equal to 1.


Write the first five terms of the sequence

Write the 55th term


1
Expert's answer
2021-02-23T10:29:01-0500
  1. The sequence has the following form: -4,4,-4,-4,4,...
  2. The sequence is the following: 87,82,77, 72, 67.
  3. The explicit formula is: "s_n=7.50", "n=1,2,3,4"; "s_n=7.50+n" for "n\geq4".

The recursive formula is: "s_n=s_{n-1}" for "n=1,2,3,4" and "s_n=s_{n-1}+1" for "n\geq4".

4.The first five terms are: 0, 3, 8, 15, 24. 55th term is: "t_n=55^2-1=3024".